OpenStudy (anonymous):

What is the first term of the geometric sequence whose fifth term is 1/24 and tenth term is 1/768 ?

OpenStudy (dan815):

A geometric sequence is where you keep multiplying by a certain ratio lets suppose the "r" be the ratio 1/24=5th term 1/24* r = 6th term 1/24*r*r=7th term 1/24 * r* r*r=8th term ... 1/24*r^5=10th term=1/768 we can solve for r

OpenStudy (dan815):

Then to figure out the first term, now that we know the 5th term and know the r value. You know that to go from Term 1 to Term 5, we would multiply Term 1 by r a total of 4 times. x* r^4=1/24 and we can find x, since we already found r previously. X is our first term.

OpenStudy (anonymous):

wat ?

OpenStudy (anonymous):

0_e

OpenStudy (dan815):

What?

OpenStudy (anonymous):

i have no clue what u typed or what your talking about _-_

OpenStudy (dan815):

Maybe it's best you did some review before attempting to solve these problems then.

OpenStudy (dan815):

Would you like me to refer you to some content/teaching material, that would be helpful to you?

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

?

OpenStudy (dan815):

What did you find your "r" to be

OpenStudy (dan815):

1/24*r^5=10th term=1/768 1/24*r^4=1/768 <--- solve for r first.

OpenStudy (anonymous):

fifth term is 1/24 and tenth is 1/768 r^5 = (1/768) / ( 1/24) = 1/32 r = 1/2 a*r^4 = 1/24 the first term: 2/3

OpenStudy (yanasidlinskiy):

Correct! You got it! It is indeed 2/3.

OpenStudy (yanasidlinskiy):

The\(\ n ^{th}\ term\ is\ ar ^{n-1}\) So we can set up two equations as follows: \(\frac{1}{24}=ar ^{4}............\)(1) \(\frac{1}{768}=ar ^{9}...........\)(2) . Hence you may divide to obtain \(r^{5}=ar ^{9}/ar ^{4}=\frac{ \frac{ 1 }{ 768 } }{ \frac{ 1 }{ 24 } }= \frac{ 1}{ 32 }\) which implies that r=1/2. Go back to the equation for 1/24 and solve to obtain a=2/3.